A Swimmer can swim at a speed of 2.8 m/s for 6 minutes. What distance will she cover in that time?
Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

swimmer rate of speed = 2.8 meters per second.

swimmer swims for 6 minutes.

6 minutes * 60 second per minute = 360 seconds.

swimmer will swim 2.8 * 360 = 1008 meters in 6 minutes.

you could also solve it by translating the speed from meters per second to meters per minute.

2.8 meters per second * 60 seconds per minute = 2.8 * 60 = 168 meters per minute.

168 meters per minute * 6 minutes = 1008 meters.


Answer by Alan3354(69443)

A Swimmer can swim at a speed of 2.8 m/s for 6 minutes. What distance will she cover in that time?
----------------
1 minute = 60 seconds.
d = r*t
--------
It's arithmetic.